OpenStudy (anonymous):

Complete the table below to solve the equation 2.5x – 10.5 = 64(0.5x). x f(x) = 2.5x – 10.5 g(x) = 64(0.5x) 2 3 4 5 6

OpenStudy (anonymous):

This is English...

OpenStudy (anonymous):

OH! my bad

OpenStudy (anonymous):

It happens :)

OpenStudy (anonymous):

But don't 4get to close it though @blueSurrency

OpenStudy (anonymous):

How...?

OpenStudy (anonymous):

:) I am a math God

OpenStudy (anonymous):

oh do you wanna help me then!

OpenStudy (anonymous):

:P not a god just really good at it.

OpenStudy (anonymous):

Well thats good too!

OpenStudy (anonymous):

Ya this is a little over my math skills but ask the following 4 help @ganeshie8 or @paki they are both really smart and can help you though...

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

this is hard xD

OpenStudy (anonymous):

i have did this befor just super hard.

OpenStudy (anonymous):

Well your help would be wonderful

OpenStudy (anonymous):

@TheSmartOne

OpenStudy (anonymous):

what are the answers?

OpenStudy (anonymous):

Its not multiple choice

OpenStudy (anonymous):

._.

OpenStudy (anonymous):

this is super hard then xD

OpenStudy (anonymous):

wait like 5 min and i have the answer xD

OpenStudy (anonymous):

That would be my exact reason for posting it! XD

OpenStudy (paki):

put the respective values of "x" in both equations separately.. .and put them in table... like for x=2.... f(2) = 2.5(2) – 10.5 and g(2) = 64(0.5(2)) f(2) = ?? and g(2) = ??

OpenStudy (anonymous):

Yeah that makes zero sense

OpenStudy (anonymous):

2.5x – 10.5 = 64(0.5x) 2.5x - 10.5 = 32x 2.5x - 32x = 10.5 -29.5x = 10.5 x = -10.5/29.5

OpenStudy (paki):

@blueSurrency http://prntscr.com/5dhhzx ????????????

OpenStudy (anonymous):

answer 47.

OpenStudy (anonymous):

But i need answers for all of it

OpenStudy (anonymous):

i think xD

OpenStudy (paki):

finding answers is your work.... sorry...

OpenStudy (anonymous):

this might help http://openstudy.com/study#/updates/531609bde4b091a77695a0b6

OpenStudy (anonymous):

I need it for 2 3 4 5 6 XD

OpenStudy (anonymous):

:P sorry best i can do.

OpenStudy (paki):

@blueSurrency "OpenStudy values the Learning process - not the ‘Give you an answer’ process"

OpenStudy (anonymous):

:p true

OpenStudy (anonymous):

It's time to sharpen the pencil and do some arithmetic! x f(x) = 2.5x - 10.5 g(x) = 64(0.5x) 2 f(x) = 2.5(2)-10.5 g(x) = 64(0.5*2) f(x) = 5-10.5 = -5.5 g(x) = 64(1) = 64 now repeat for x = 3, 4, 5, 6 was from whpalmer4 so yeh he is also a nother 99 so i have to say xD

OpenStudy (anonymous):

Okay im loving the sarcasm but im just trying to answer this friggin question

OpenStudy (anonymous):

:P

OpenStudy (anonymous):

you know it is sad when a mathlet and a human calculator cant do this./

OpenStudy (anonymous):

No no sweetie you have been a huge help! Thank you so much!

OpenStudy (anonymous):

but if you post this in the math section they might beable to help you better.

OpenStudy (anonymous):

I know im new to open study so i didnt know there were different sections but thank you(:

OpenStudy (anonymous):

well welcome to os :)

OpenStudy (anonymous):

if you ever need anyone's help just do @Name. and click their name when it pops up.

OpenStudy (anonymous):

or if you have any questions you can go to the os feedback section and there they will answer your questions.

OpenStudy (anonymous):

Okie thank you(:

OpenStudy (anonymous):

any time.